# RadarGaussianDet3D: An Efficient and Effective Gaussian-based 3D Detector with 4D Automotive Radars

### v1

@@ Abstract
4D automotive radars have gained increasing attention for autonomous driving due to their low cost, robustness, and inherent velocity measurement capability. However, existing 4D radar-based 3D detectors rely heavily on pillar encoders for BEV feature extraction, where each point contributes to only a single BEV grid, resulting in sparse feature maps and degraded representation quality. In addition, they also optimize bounding box attributes independently, leading to sub-optimal detection accuracy. Moreover, their inference speed, while sufficient for high-end GPUs, may fail to meet the real-time requirement on vehicle-mounted embedded devices. To overcome these limitations, an efficient and effective Gaussian-based 3D detector, namely RadarGaussianDet3D is introduced, leveraging Gaussian primitives and distributions as intermediate representations for radar points and bounding boxes. In RadarGaussianDet3D, a novel Point Gaussian Encoder (PGE) is designed to transform each point into a Gaussian primitive after feature aggregation and employs the 3D Gaussian Splatting (3DGS) technique for BEV rasterization, yielding denser feature maps. PGE exhibits exceptionally low latency, owing to the optimized algorithm for point feature aggregation and fast rendering of 3DGS. In addition, a new Box Gaussian Loss (BGL) is proposed, which converts bounding boxes into 3D Gaussian distributions and measures their distance to enable more comprehensive and consistent optimization. Extensive experiments on TJ4DRadSet and View-of-Delft demonstrate that RadarGaussianDet3D achieves state-of-the-art detection accuracy while delivering substantially faster inference, highlighting its potential for real-time deployment in autonomous driving.
@@ I Introduction
Within the perception framework, accurate and robust 3D object detection serves as a cornerstone for safe navigation and decision-making in autonomous driving. This paper addresses three major challenges encountered by existing 4D radar-based 3D detectors. First, most detectors adopt the pillar encoder in PointPillars [3] to construct radar bird’s-eye-view (BEV) feature maps. As noted by Lin et al. [4] and shown in Fig. 1(a), the scatter operation in the pillar encoder maps sparse radar points to individual BEV grids based on their coordinates, producing sparse feature maps. Although stacking additional BEV encoder layers can mitigate this problem, small object features may blend into background regions. RCBEVDet [4] proposes an enhanced scatter operation that assigns each point feature to all neighboring grids (Fig. 1(b)), with neighborhood size determined by the radar cross section (RCS), which roughly reflects object size. However, this hand-crafted design is sub-optimal, since RCS is influenced not only by object size but also by factors such as material properties. In addition, if a grid lies within the neighborhoods of multiple points, only the maximum feature value is retained, leading to information loss. Therefore, a more suitable feature scattering strategy must be explored.
@@
To address these issues, RadarGaussianDet3D is proposed as an efficient and effective 4D radar-based 3D object detector that employs Gaussian primitives and distributions as intermediate representations of radar points and bounding boxes. For the first problem, the Point Gaussian Encoder (PGE) is designed to generate BEV feature maps using the 3D Gaussian Splatting (3DGS) technique [6]. 3DGS represents a scene with 3D Gaussian primitives, where a view is synthesized by projecting them onto a plane followed by differentiable rasterization. Since a 3D Gaussian primitive can be regarded as a point with additional attributes such as size and rotation, it is natural to predict these attributes from radar points. The BEV feature map is then obtained through Gaussian splatting, which acts as a feature scattering strategy (Fig. 1(c)). In this process, each point contributes to all grids overlapped by the projected 2D Gaussian primitive, and alpha-blending is used to integrate features from multiple points. For bounding box regression, a Box Gaussian Loss (BGL) is proposed, which transforms predicted and ground-truth boxes into 3D Gaussian distributions and computes their distance, thereby comprehensively optimizing all box attributes. Finally, since 3DGS [6] provides ultra-fast rendering and BGL is used only during training, the proposed model achieves high inference efficiency.
@@
  • A 4D radar-based 3D object detection model, RadarGaussianDet3D, is proposed. In this model, the Point Gaussian Encoder (PGE) replaces conventional voxel or pillar encoders to generate BEV feature maps, while the Box Gaussian Loss (BGL) is introduced to assist learning.
  • In PGE, radar points are represented as 3D Gaussian primitives. Point features are extracted by the Local Feature Aggregation (LFA) and Global Feature Aggregation (GFA) modules in parallel, from which Gaussian attributes are predicted. BEV Gaussian splatting is then employed to produce denser BEV feature maps, alleviating the impact of point sparsity.
  • For BGL, both ground-truth and predicted bounding boxes are converted into 3D Gaussian distributions based on their positions, sizes, and orientations. The distance between distributions is computed as loss, enabling comprehensive optimization of box attributes.
  • Experiments on TJ4DRadSet [7] and View-of-Delft [8] demonstrate that RadarGaussianDet3D achieves substantially faster inference while maintaining accuracy comparable to state-of-the-art methods. Ablation studies further validate the effectiveness of each proposed module.
The remainder of this paper is organized as follows. Section II reviews related work, including 4D radar-based 3D object detection methods and applications of 3D Gaussian splatting in autonomous driving. Section III presents the proposed RadarGaussianDet3D model in detail, and Section IV reports and analyzes the experimental results. Finally, Section V concludes the paper.
@@ II Related Work
Several approaches focus on enriching radar feature representations [9, 10, 5, 11]. For instance, RCFusion [10] processes spatial, velocity, and intensity values separately to avoid feature confusion. RadarPillars [5] introduces self-attention among non-empty pillars to aggregate features belonging to the same object. SMURF [11] predicts point density distributions using kernel density estimation as an additional feature, improving sparsity-awareness in detection.
@@
As certain limitations of 4D radar cannot be fully resolved algorithmically, some methods leverage auxiliary modalities during training. For example, SCKD [13] employs cross-modality distillation to transfer knowledge from LiDAR to 4D radar, enhancing performance without compromising efficiency. Other approaches explore multi-modal fusion, integrating semantic cues from RGB images [14, 15, 16, 17] or geometric information from LiDAR point clouds [18, 19, 20] to complement 4D radar.
@@
3D Gaussian Splatting (3DGS) [6] introduced a new paradigm in 3D reconstruction. Although initially proposed for camera-based novel view synthesis (NVS) in static scenes, researchers have extended it to other domains, including autonomous driving. For instance, SplatAD [21] and RadarSplat [2] extend 3DGS to NVS with LiDAR and radar sensors, respectively, while DrivingGaussian [22] adapts it to dynamic scenes.
3DGS has also been applied to perception tasks such as BEV segmentation [23, 24], 3D object detection [25], and occupancy prediction [26, 27, 28]. However, perception tasks require strong generalization ability and real-time performance, making per-scene optimization of Gaussian attributes as in the original implementation infeasible. Consequently, perception methods predict Gaussian attributes in a forward pass, either from feature maps [23, 24, 29] or from learnable queries [26, 27]. The former is simple and direct, with each Gaussian primitive predicted from a feature map element (e.g., an image pixel), making it suitable for single-modality inputs. In contrast, the latter is more flexible, allowing Gaussian primitives to be generated in varying quantities and supporting multi-modal inputs. Some works even extend this query-based strategy to end-to-end autonomous driving [30, 31] and world models [28].
@@ III-A Overview
III-B Point Gaussian Encoder (PGE): Radar Points as Gaussian Primitives
@@
where $p_{i}\in\mathbb{R}^{3},f_{i}\in\mathbb{R}^{C_{\textup{raw}}}$ denote the 3D coordinates and raw features of the $i$-th point $(i=1,\cdots,N)$, respectively. Here, $f^{i}_{\textup{LFA}}\in\mathbb{R}^{C}$ represents the updated features after LFA. $N$ denotes the total number of points, and $C$ refers to the channel dimension. The $i$-th point is referred to as the center point, while points in $\mathcal{N}_{i}$ are its neighbors. In addition, the position offset from the center point to each neighbor is appended to the features, following the pillar feature extraction in PointPillars [3].
To avoid traversing all points for averaging neighbor features and to reduce memory usage, an optimized algorithm is devised. Its core relies on the PyTorch-supported operation $\mathtt{scatter\_reduce}$, as illustrated in Algorithm 1 and visualized in Fig. 3(c). Experiments in Section IV-C confirm that the proposed Indexing & Scattering algorithm achieves significantly higher speed than the Traversal method in Fig. 3(a), while requiring considerably less memory compared with the Broadcasting & Masking strategy in Fig. 3(b).
Algorithm 1 Local Feature Aggregation (Indexing & Scattering).
0:  Point coordinates $p\in\mathbb{R}^{N\times 3}$, point features $f\in\mathbb{R}^{N\times{C_{\textup{raw}}}}$, radius $r$
0:  Updated point features $f_{\textup{LFA}}\in\mathbb{R}^{N\times C}$
1:  $D\leftarrow\mathtt{L2\_norm}(p.\mathtt{unsqueeze}(0)-p.\mathtt{unsqueeze}(1),\textup{dim=2})\triangleright$ Pair-wise distance, $N\times N$
2:  $M\leftarrow(D<r)\triangleright$ Mask, $N\times N$
3:  $i_{\textup{row}},j_{\textup{column}}\leftarrow\mathtt{nonzero}(M)\triangleright$ Find non-zero positions
4:  $p_{\textup{center}}\leftarrow p[i_{\textup{row}}]\triangleright$ Center point locations
5:  $p_{\textup{neighbor}}\leftarrow p[j_{\textup{column}}]\triangleright$ Neighbor point positions
6:  $f_{\textup{neighbor}}\leftarrow f[j_{\textup{column}}]\triangleright$ Neighbor point features
7:  $f_{\textup{neighbor}}\leftarrow\mathtt{Concat}([f_{\textup{neighbor}},p_{\textup{center}}-p_{\textup{neighbor}}])\triangleright$ Append neighbor point features with positionoffsets from center points to neighbor points
8:  $f_{\textup{neighbor}}\leftarrow\mathtt{Linear}(f_{\textup{neighbor}})\triangleright$ Feature projection
9:  $f_{\textup{LFA}}\leftarrow\mathtt{scatter\_reduce}(\textup{dim=0},\textup{index=}i_{\textup{row}},\textup{source=}f_{\textup{neighbor}},\newline \textup{\hskip 99.58464ptreduce=`mean'})$ $\triangleright$ Average all neighbor point features for each center point
10:  return $f_{\textup{LFA}}$
@@
where $f\in\mathbb{R}^{N\times{C_{\textup{raw}}}}$ denotes the raw point features, and $f_{\textup{GFA}}\in\mathbb{R}^{N\times C}$ represents the features after GFA. Here, $\mathtt{SelfAttn}(Q,K,V)$ is the self-attention operation with queries $Q$, keys $K$, and values $V$, while $\mathtt{FFN}(\cdot)$ denotes the feed-forward network.
@@
$s,q,f^{g}=\mathtt{Linear}(\mathtt{Concat}([f,f_{\textup{LFA}},f_{\textup{GFA}}])),$ (3)
where $q$ is normalized to satisfy the unit-norm constraint of quaternions.
@@
III-C Box Gaussian Loss (BGL): Bounding Boxes as Gaussian Distributions
@@
Box-to-Gaussian Conversion. Since 3D bounding boxes and 3D Gaussian distributions share attributes such as position, scale, and orientation, they can be mutually converted. Each ground-truth bounding box $b=[x,y,z,l,w,h,\theta]$ is transformed into a 3D Gaussian distribution $g=\mathcal{N}(\mu;\mathbf{\Sigma})$ as follows:
@@
where $N_{b}$ denotes the number of ground-truth bounding boxes, and
represents the KL divergence between the predicted Gaussian distribution $\hat{g}$ and the ground-truth one $g$.
The first term $(\hat{\mu}-\mu)^{\top}\mathbf{\Sigma}^{-1}(\hat{\mu}-\mu)$ is the Mahalanobis distance between $\hat{\mu}$ and $\mathcal{N}(\mu;\mathbf{\Sigma})$. Unlike the $L_{1}$ distance, it accounts for the shape and orientation of the ground-truth object, thereby providing a more informative localization error. The remaining terms further capture differences in scales and orientations.
The total regression loss $\mathcal{L}_{\text{reg}}$ is defined as
@@ IV-A Implementation Details
Datasets and Evaluation Metrics. Two publicly available datasets are used to evaluate the proposed RadarGaussianDet3D: TJ4DRadSet [7] and View-of-Delft (VoD) [8]. Both datasets provide synchronized data from a 4D radar, a camera, and a LiDAR, with tens of thousands of annotated 3D bounding boxes. Compared with VoD, TJ4DRadSet is collected under more complex road and traffic conditions with greater lighting variability, making detection significantly more challenging.
@@
    Inference speeds are measured on Tesla V100, RTX 3090, RTX 4090 (mainstream GPUs) or AGX Xavier (an embedded device designed for autonomous driving).
@@
Network and Hyper-parameters. The proposed model is developed based on CenterPoint [33], with the pillar encoder replaced by the proposed PGE and the BGL incorporated. In BEV Gaussian splatting, the target feature map size $(H,W)$ is set to $(320,320)$ for VoD and $(432,496)$ for TJ4DRadSet, ensuring the same resolution as the output of the pillar encoder with the official pillar size of $0.16\mathrm{m}$. The spherical neighborhood radius in LFA is set to $r=0.32\mathrm{m}$, and the feature dimension in PGE is fixed to $C=64$.

### v2

# RadarGaussianDet3D: Gaussian Representation-based Real-time 3D Object Detection with 4D Automotive Radars

@@ Abstract
4D automotive radars have gained increasing attention for autonomous driving due to their low cost, robustness, and inherent velocity measurement capability. However, existing 4D radar-based 3D detectors rely heavily on pillar encoders for BEV feature extraction, where each point contributes to only a single BEV grid, resulting in sparse feature maps and degraded representation quality. In addition, they also optimize bounding box attributes independently, leading to sub-optimal detection accuracy. Moreover, their inference speed, while sufficient for high-end GPUs, may fail to meet the real-time requirement on vehicle-mounted embedded devices. To overcome these limitations, an efficient and effective Gaussian-based 3D detector, namely RadarGaussianDet3D is introduced, leveraging Gaussian primitives and distributions as intermediate representations for radar points and bounding boxes. In RadarGaussianDet3D, a novel Point Gaussian Encoder (PGE) is designed to transform each point into a Gaussian primitive after feature aggregation and employs the 3D Gaussian Splatting (3DGS) technique for BEV rasterization, yielding denser feature maps. PGE exhibits exceptionally low latency, owing to the optimized algorithm for point feature aggregation and fast rendering of 3DGS. In addition, a new Box Gaussian Loss (BGL) is proposed, which converts bounding boxes into 3D Gaussian distributions and measures their distance to enable more comprehensive and consistent optimization. Extensive experiments on TJ4DRadSet and View-of-Delft demonstrate that RadarGaussianDet3D achieves high detection accuracy while delivering substantially faster inference, highlighting its potential for real-time deployment in autonomous driving. Source code is available at https://github.com/XiongWeiyi/RadarGaussianDet3D.
@@ I Introduction
Within the perception framework, accurate and robust 3D object detection serves as a cornerstone for the subsequent trajectory prediction and motion planning in autonomous driving. This paper addresses three major challenges encountered by existing 4D radar-based 3D detectors. First, most detectors adopt the pillar encoder in PointPillars [14] to construct radar bird’s-eye-view (BEV) feature maps. As noted by Lin et al. [16] and shown in Fig. 1(a), the scatter operation in the pillar encoder maps sparse radar points to individual BEV grids based on their coordinates, producing sparse feature maps. Although stacking additional BEV encoder layers can mitigate this problem, small object features may blend into background regions. RCBEVDet [16] proposes an enhanced scatter operation that assigns each point feature to all neighboring grids (Fig. 1(b)), with neighborhood size determined by the radar cross section (RCS), which roughly reflects object size. However, this hand-crafted design is sub-optimal, since RCS is influenced not only by object size but also by factors such as material properties. In addition, if a grid lies within the neighborhoods of multiple points, only the maximum feature value is retained, leading to information loss. Therefore, a more suitable feature scattering strategy must be explored.
@@
To address these issues, RadarGaussianDet3D is proposed as an efficient and effective 4D radar-based 3D object detector that employs Gaussian primitives and distributions as intermediate representations of radar points and bounding boxes. For the first problem, the Point Gaussian Encoder (PGE) is designed to generate BEV feature maps using the 3D Gaussian Splatting (3DGS) technique [12]. 3DGS represents a scene with 3D Gaussian primitives, where a view is synthesized by projecting them onto a plane followed by differentiable rasterization. Since a 3D Gaussian primitive can be regarded as a point with additional attributes such as size and rotation, it is natural to predict these attributes from radar points. Note that, unlike the original 3DGS framework that requires per-scene optimization, this process is feed-forward, single-shot, and generalizable. The BEV feature map is then obtained through Gaussian splatting, which acts as a feature scattering strategy (Fig. 1(c)). In this process, each point contributes to all grids overlapped by the projected 2D Gaussian primitive, and alpha-blending is used to integrate features from multiple points. For bounding box regression, a Box Gaussian Loss (BGL) is proposed, which transforms predicted and ground-truth boxes into 3D Gaussian distributions and computes their distance, thereby comprehensively optimizing all box attributes. Finally, since 3DGS [12] provides ultra-fast rendering and BGL is used only during training, the proposed model achieves high inference efficiency.
@@
  • Radar points are modeled as 3D Gaussian primitives, and a Point Gaussian Encoder (PGE) is designed to efficiently splat these primitives into BEV feature maps, effectively alleviating the inherent sparsity of radar point clouds.
  • 3D bounding boxes are formulated as Gaussian distributions, and a Box Gaussian Loss (BGL) is introduced to measure box discrepancies via distributional distance, enabling joint optimization of all bounding box attributes.
  • By integrating PGE and BGL, a 4D radar-based 3D object detection framework, termed RadarGaussianDet3D, is developed. Experiments on the TJ4DRadSet[36] and View-of-Delft[21] datasets demonstrate a favorable accuracy–efficiency trade-off, achieving substantially lower latency than state-of-the-art methods while maintaining high detection accuracy.
This paper is organized as follows. Section II reviews related work, including 4D radar-based 3D object detection methods and applications of 3D Gaussian splatting in autonomous driving. Section III presents the proposed RadarGaussianDet3D model in detail, and Section IV reports and analyzes the experimental results. Finally, Section V concludes the paper.
@@ II Related Work
Several approaches focus on enriching radar features [29, 34, 20, 17]. For instance, RCFusion [34] processes spatial, velocity, and intensity values separately to avoid feature confusion. RadarPillars [20] introduces self-attention among non-empty pillars to aggregate features belonging to the same object. SMURF [17] predicts point density distributions using kernel density estimation as an additional feature, improving sparsity-awareness in detection.
@@
As certain limitations of 4D radar cannot be fully resolved algorithmically, some methods use auxiliary modalities during training. E.g., SCKD [30] employs cross-modality distillation to transfer knowledge from LiDAR to 4D radar, enhancing performance without compromising efficiency. Other approaches explore multi-modal fusion, integrating geometric information from LiDAR point clouds [5, 9, 25] or semantic cues from RGB images [27, 28, 1, 35] to complement 4D radar.
In contrast to the aforementioned methods, this work addresses the sparsity of 4D radar point clouds at the representation level. Instead of relying on the commonly used pillar-based representation, radar points are modeled as 3D Gaussian primitives, and these primitives are rasterized using 3DGS to generate denser BEV feature maps.
@@
3D Gaussian Splatting (3DGS) [12] introduced a new paradigm in 3D reconstruction. Although initially proposed for camera-based novel view synthesis (NVS) in static scenes, researchers extended it to other domains, including autonomous driving. For instance, SplatAD [8] and RadarSplat [13] extend 3DGS to NVS with LiDAR and radar sensors, respectively, while DrivingGaussian [38] adapts it to dynamic scenes.
3DGS has also been applied to various perception tasks, including BEV segmentation [4, 19], 3D object detection [2], and occupancy prediction [11, 10, 39]. However, perception tasks typically require strong generalization capability and real-time inference, making per-scene optimization of Gaussian attributes in the original 3DGS formulation impractical. As a result, existing perception-oriented methods predict Gaussian attributes in a feed-forward manner, either directly from feature maps [4, 19, 31] or via learnable queries [11, 10]. In the feature-based paradigm, each Gaussian primitive is predicted directly from a feature map element (e.g., an image pixel), resulting in a simple and efficient design that is particularly suitable for single-modality inputs. Moreover, this one-shot prediction incurs minimal computational overhead, making it highly favorable for real-time systems. In contrast, query-based methods rely on iterative refinement of Gaussian primitives through repeated query-data interactions. While this design allows a variable number of Gaussian primitives and naturally supports multi-modal inputs, it introduces additional computational cost. Beyond perception subtasks, query-based Gaussian representations have also been extended to end-to-end autonomous driving [18, 37] and world models [39].
@@
While existing approaches have achieved promising results, most focus on image-view transformation or multi-modal feature fusion. In contrast, this work pioneers the application of 3DGS to 4D radar point clouds for BEV feature extraction, where Gaussian attributes are directly predicted from sparse radar measurements. This design establishes a new paradigm for leveraging 3DGS in radar-based perception tasks.
@@ III-A Overview
III-B Point Gaussian Encoder (PGE)
@@
where $p_{i}\in\mathbb{R}^{3}$ and $f_{i}\in\mathbb{R}^{C_{\textup{raw}}}$ denote the 3D coordinates and raw features of the $i$-th point $(i=1,\cdots,N)$, respectively. Here, $f^{i}_{\textup{LFA}}\in\mathbb{R}^{C}$ represents the updated features after LFA. $N$ denotes the total number of points, and $C$ refers to the channel dimension. The $i$-th point is referred as the center point, while points in $\mathcal{N}_{i}$ are its neighbors. In addition, the position offset from the center point to each neighbor is appended to features, following the pillar feature extraction in PointPillars [14].
To avoid traversing all points for averaging neighbor features and to reduce memory usage, an optimized algorithm is devised. It relies on the PyTorch-supported operation $\mathtt{scatter\_reduce}$, as visualized in Fig. 3(c). Experiments in Section IV-C confirm that the proposed Indexing & Scattering algorithm achieves higher speed than the Traversal method in Fig. 3(a), while requiring considerably less memory compared with the Broadcasting & Masking strategy in Fig. 3(b).
@@
where $f\in\mathbb{R}^{N\times{C_{\textup{raw}}}}$ denotes the raw point features, and $f_{\textup{GFA}}\in\mathbb{R}^{N\times C}$ represents the features after GFA. Here, $\mathtt{SelfAttn}(Q,K,V)$ is the self-attention operation [24] with queries $Q$, keys $K$, and values $V$, while $\mathtt{FFN}(\cdot)$ denotes the feed-forward network.
@@
$s^{\prime},q,f^{g}$ $=\mathtt{Linear}(\mathtt{Concat}([f,f_{\textup{LFA}},f_{\textup{GFA}}])),$ (3)
$s$ $=\mathtt{Sigmoid}(s^{\prime})\cdot s_{\max},$
where $\mathtt{Sigmoid}$ ensures positive scale values while preventing excessive magnitude following[33], and $s_{\max}=1\,\mathrm{m}$ is a predefined maximum scale. $q$ is normalized to satisfy the unit-norm constraint of quaternions.
Compared with original 3DGS paradigm, where Gaussian attributes are treated as learnable parameters optimized per scene, the proposed feed-forward prediction strategy ensures model generalizability across scenes. Moreover, the single-shot prediction is computationally more efficient than the iterative refinement in query-based Gaussian prediction methods.
@@
III-C Box Gaussian Loss (BGL)
@@
Box-to-Gaussian Conversion. Since 3D bounding boxes and 3D Gaussian distributions share attributes ( e.g., position, scale, and orientation), they can be mutually converted. Each ground-truth bounding box $b=[x,y,z,l,w,h,\theta]$ is transformed into 3D Gaussian distribution $g=\mathcal{N}(\mu;\mathbf{\Sigma})$ as follows:
@@
where $N_{b}$ denotes the number of ground-truth bounding boxes;
[…]
The first term $(\hat{\mu}-\mu)^{\top}\mathbf{\Sigma}^{-1}(\hat{\mu}-\mu)$ is the Mahalanobis distance between $\hat{\mu}$ and $\mathcal{N}(\mu;\mathbf{\Sigma})$. Unlike $L_{1}$ distance, it accounts for the shape and orientation of the ground-truth object, thereby providing more informative localization error. The remaining terms capture differences in scales and orientations.
The total regression loss $\mathcal{L}_{\text{reg}}$ is defined as
@@ IV-A Implementation Details
Datasets and Evaluation Metrics. Two publicly available datasets are used to evaluate the proposed RadarGaussianDet3D: TJ4DRadSet [36] and View-of-Delft (VoD) [21]. Both datasets provide synchronized data from a 4D radar, a camera, and a LiDAR, with tens of thousands of annotated 3D bounding boxes. Compared with VoD, TJ4DRadSet is collected under more complex road and traffic conditions with greater lighting variability, making detection significantly more challenging. Experiments on both TJ4DRadSet and VoD are conducted using their official dataset partitioning, with models trained and evaluated on each dataset independently.
@@
    Inference speeds are measured on Tesla V100, RTX 3090, RTX 4090 or AGX Xavier (an embedded device designed for autonomous driving).
@@
Network and Hyper-parameters. The proposed model is developed based on CenterPoint-Pillar [32], with the pillar encoder replaced by the proposed PGE and the BGL incorporated. In BEV Gaussian splatting, the target feature map size $(H,W)$ is set to $(320,320)$ for VoD and $(432,496)$ for TJ4DRadSet, ensuring the same resolution as the output of the pillar encoder with the commonly adopted pillar size of $0.16\mathrm{m}$ [17, 20, 3], for fair comparison. The spherical neighborhood radius in LFA is set to $r=0.32\mathrm{m}$, where small deviations have only a marginal impact on detection accuracy, and the feature dimension in PGE is fixed to $C=64$.